if an equation is 0.249=(2.61/2.53X) How do you find X
Answer by stanbon(75887)

if an equation is 0.249=(2.61/2.53X) How do you find X
-------
Invert both sides to get:
2.53x/2.61 = 1/0.249
-------------------------------
Multiply both sides by 2.61
Then divide both sides by 2.53 to get:
x = (2.61/2.53)(1/0.249)
----
x = 4.1431
============================
Cheers,
Stan H.